# Learnable Sparsity for Vision Generative Models

### v2

@@ 7 Conclusion
This work proposes EcoDiff, a structural pruning framework for large vision generative models that significantly reduces computational and memory requirements without sacrificing generation quality. By enabling efficient pruning and lightweight post-pruning retraining, our approach lowers the barrier to deploying state-of-the-art text-to-image systems on more modest hardware, promoting broader accessibility of generative AI technology. In addition, reducing model size and retraining cost can contribute to lower energy consumption and carbon footprint, helping mitigate the environmental impact of large-scale model training. We recognize that generative models may still pose societal risks, such as bias or misuse, and encourage responsible deployment of pruned models within appropriate ethical and regulatory frameworks.
We also disclose that we used large language models only in a supportive role for polishing and formatting. All technical contributions, decisions, and claims are the authors’ own.
@@
All source code, scripts, and configuration files required to reproduce our experiments are included in the repository. We make the full, documented repository publicly available on GitHub. A detailed description of the datasets and hyperparameters can be found in the appendix.
@@ Appendix A Derivation of the Differentiable $L_{0}$ Regularization Loss
In this section, we provide the detailed derivation of the expected $L_{0}$ complexity loss $\mathcal{L}_{0}(\bm{\lambda})$ presented in Equation 12. We adopt the hard-concrete sampling mechanism proposed by (louizos2018learning).
@@ A.2 Deriving the $L_{0}$ Loss
In our experiments, we adopt the standard hyperparameter settings established in prior work. We set the stretch parameters to $\gamma=-0.1$ and $\zeta=1.1$ following louizos2018learning; xia2024sheared. For the temperature parameter, we use $\alpha=0.83$ following the configuration in Sheared LLaMA (xia2024sheared).
@@ Appendix B Detailed Experiment Setting and Evaluation Setting
Model: we prune SOTA latent diffusion models, SDXL, and FLUX, representing various latent diffusion architectures (rombach2022high; podell2024sdxl; blackforestlabs2024). For FLUX, we employ the time-step distillation model, FLUX.1-schnell. Pruning details: we only use text prompts from GCC3M (sharma2018conceptual) for training. Baselines: we compare our method with the SOTA diffusion pruning methods, DiffPruning (fang2023structural) and BK-SDM (kim2024bk). For BK-SDM, we only run on SD-XL, since the method is constrained for UNet models. To ensure a fair evaluation, we perform the pruning under the same computation budget of 10 A100 GPU hours and 100 data samples. Evaluation metrics: we select FID (heusel2017gans), CLIP score (radford2021learning), and SSIM (wang2004image). We evaluate SSIM by comparing the pruned and original (unmasked) models. We use pretrained CLIP encoder ViT-B-16 to calculate the CLIP score. For quantitative evaluation, we use the MS COCO and Flickr 30k datasets (lin2014microsoft; young2014image), randomly selecting 5,000 image-caption pairs from each. Additionally, we measure computation in Giga Floating Point Operation (GFLOP) and the total number of parameters in the pruned models. Miscellaneous: To ensure reproducibility, we use the Hugging Face Diffusers and Accelerator library (accelerate; diffusers). All experiments are conducted on a single NVIDIA A100 GPU with 80G VRAM.
@@ E.2 Compatibility With Feature Reuse
EcoDiff demonstrates high versatility by efficiently integrating with feature reuse methods like DeepCache (ma2024deepcache) as shown in Figure 6. Additionally, combining EcoDiff with DeepCache enables significant reductions in effective GFLOPs and runtime, highlighting its compatibility and effectiveness in optimizing resource usage.
@@ F.1 Details of Post-Pruning Adaptation
To mitigate this degradation, we adopt two complementary retraining strategies: LoRA adaptation and full-model fine-tuning. In both cases, retraining is performed for 10,000 steps and requires approximately 10 hours of additional compute. Importantly, the training images are not sourced from the original dataset but are instead synthesized by the original SDXL model using prompts from GCC3M (sharma2018conceptual).
@@ F.2 Post-Pruning Retraining Ablation on Datasets
To address potential concerns that this approach might limit the validity of our recovery claims, we conducted an additional ablation study using a public subset of the LAION2B-en-aesthetic dataset (schuhmann2022laion5bopenlargescaledataset), derived from the LAION-5B corpus. This subset is generally considered to share a distribution closer to the models’ original training data than synthetic imagery.
@@ Appendix I Carbon Footprint Analysis
We conduct a carbon footprint analysis for EcoDiff training. The analysis uses training configurations with only 200 training steps, as most configurations converge within 400 steps. The carbon footprint calculations are based on the default configurations outlined in Appendix B. All calculations follow the methodology used in the SD2 carbon footprint analysis (rombach2022high).

### v3

@@ 7 Conclusion
Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA2386-24-1-4011, and this research is partially supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2509).
@@ Acknowledgments
This work proposes EcoDiff, a structural pruning framework for large vision generative models that substantially reduces computational and memory costs while preserving generation quality. By enabling efficient pruning with lightweight post-pruning retraining, EcoDiff lowers the hardware and energy requirements for deploying state-of-the-art text-to-image systems, improving accessibility and reducing environmental impact. As with all generative models, ethical risks such as bias and misuse remain; we therefore emphasize responsible deployment of pruned models in accordance with established ethical guidelines and regulatory standards.
LLM Usage Statement
We disclose that we used large language models (LLMs) only in a supportive role for polishing and formatting. All technical contributions, decisions, and claims are the authors’ own.
@@
All source code, scripts, and configuration files required to reproduce our experiments are included in the repository. We make the full, documented repository publicly available on GitHub. A detailed description of the datasets and hyper-parameters can be found in the appendix B.
@@ Appendix A Derivation of the Differentiable $L_{0}$ Regularization Loss
In this section, we provide the detailed derivation of the expected $L_{0}$ complexity loss $\mathcal{L}_{0}(\bm{\lambda})$ presented in Equation 12. We adopt the hard-concrete sampling mechanism proposed by (Louizos et al., 2018).
@@ A.2 Deriving the $L_{0}$ Loss
In our experiments, we adopt the standard hyperparameter settings established in prior work. We set the stretch parameters to $\gamma=-0.1$ and $\zeta=1.1$ following Louizos et al. (2018); Xia et al. (2024). For the temperature parameter, we use $\alpha=0.83$ following the configuration in Sheared LLaMA (Xia et al., 2024).
@@ Appendix B Detailed Experiment Setting and Evaluation Setting
Model: we prune SOTA latent diffusion models, SDXL, and FLUX, representing various latent diffusion architectures (Rombach et al., 2022; Podell et al., 2024; Black Forest Labs, 2024). For FLUX, we employ the time-step distillation model, FLUX.1-schnell. Pruning details: we only use text prompts from GCC3M (Sharma et al., 2018) for training. Baselines: we compare our method with the SOTA diffusion pruning methods, DiffPruning (Fang et al., 2023) and BK-SDM (Kim et al., 2024). For BK-SDM, we only run on SD-XL, since the method is constrained for UNet models. To ensure a fair evaluation, we perform the pruning under the same computation budget of 10 A100 GPU hours and 100 data samples. Evaluation metrics: we select FID (Heusel et al., 2017), CLIP score (Radford et al., 2021), and SSIM (Wang et al., 2004). We evaluate SSIM by comparing the pruned and original (unmasked) models. We use pretrained CLIP encoder ViT-B-16 to calculate the CLIP score. For quantitative evaluation, we use the MS COCO and Flickr 30k datasets (Lin et al., 2014; Young et al., 2014), randomly selecting 5,000 image-caption pairs from each. Additionally, we measure computation in Giga Floating Point Operation (GFLOP) and the total number of parameters in the pruned models. Miscellaneous: To ensure reproducibility, we use the Hugging Face Diffusers and Accelerator library (Hugging Face, 2021; 2022). All experiments are conducted on a single NVIDIA A100 GPU with 80G VRAM.
@@ E.2 Compatibility With Feature Reuse
EcoDiff demonstrates high versatility by efficiently integrating with feature reuse methods like DeepCache (Ma et al., 2024b) as shown in Figure 6. Additionally, combining EcoDiff with DeepCache enables significant reductions in effective GFLOPs and runtime, highlighting its compatibility and effectiveness in optimizing resource usage.
@@ F.1 Details of Post-Pruning Adaptation
To mitigate this degradation, we adopt two complementary retraining strategies: LoRA adaptation and full-model fine-tuning. In both cases, retraining is performed for 10,000 steps and requires approximately 10 hours of additional compute. Importantly, the training images are not sourced from the original dataset but are instead synthesized by the original SDXL model using prompts from GCC3M (Sharma et al., 2018).
@@ F.2 Post-Pruning Retraining Ablation on Datasets
To address potential concerns that this approach might limit the validity of our recovery claims, we conducted an additional ablation study using a public subset of the LAION2B-en-aesthetic dataset (Schuhmann et al., 2022), derived from the LAION-5B corpus. This subset is generally considered to share a distribution closer to the models’ original training data than synthetic imagery.
@@ Appendix I Carbon Footprint Analysis
We conduct a carbon footprint analysis for EcoDiff training. The analysis uses training configurations with only 200 training steps, as most configurations converge within 400 steps. The carbon footprint calculations are based on the default configurations outlined in Appendix B. All calculations follow the methodology used in the SD2 carbon footprint analysis (Rombach et al., 2022).